OpenStudy (anonymous):

negative 3 + negative 7?

OpenStudy (anonymous):

\[-3 + ( -7 ) = -3 - 7 \]

OpenStudy (anonymous):

negative 10

OpenStudy (anonymous):

@zscdragon i think u just gave me an aneurysm tbh

OpenStudy (skullpatrol):

Nope the answer is $$\Huge-3 + (-7) = -(|-3|+|-7|)$$$$\Huge=-(10)=-10$$

OpenStudy (anonymous):

heavy breathing should i like switch the medals out and give it to skullpatrol now or